### v2

@@ 1 Introduction
We will demonstrate in this paper that one can stabilize the axion in an $SL(2,\mathbb{Z})$-compatible manner in the models based on $j$-functions in Kallosh:2024ymt . The essential idea is to employ the $SL(2,\mathbb{Z})$ invariance of the $j$-function, and hence to introduce axion stabilization via terms such as as $|j(\tau)\pm\overline{j(\tau)}|^{2}$, or more general terms, such as $|j(\tau)\,e^{-i2\pi\gamma}-\overline{j(\tau)}\,e^{i2\pi\gamma}|^{2}$ where $\gamma$ is some constant. As we will see, this leads to the formation of axion valleys, stabilizing the axion field. Inflation in such two-field models occurs in the same way as in the single-field $\alpha$-attractors, and therefore leads to the same cosmological predictions.
@@ 2 Axion stabilization preserving $SL(2,\mathbb{Z})$ invariance
In the present paper we will outline a number of possibilities to avoid this potential problem by making the axion massive and stabilizing its value during inflation. As a simple example of a model with axion stabilization, consider a potential
@@
Indeed this leads to the axion stabilization at $\theta=k/2$.
@@
Figure 2: Left panel (2(a)): A stream plot of the slow-roll inflationary trajectories for $A=10^{{-3}}$, $13\alpha=$ 1 in the range $20<\theta<1/$ 2, $65<\varphi<$ 6. Even the trajectories beginning very close to the ridge at $\theta=1/4$ rapidly reach the axion valleys at $\theta=0$ or $\theta=1/2$. This process is even more efficient for the trajectories at greater values of $\varphi$ and $A$. Right panel (2(b)): The same stream plot, for $A=10^{{-2}}$.
@@
Figure 3: The slow-roll parameter $\eta^{\theta}{}_{\theta}$ (14) at $\varphi=2$ for $13\alpha=$ 1, $\beta=12^{3}$. In the left panel $A=1/10$, in the right one $A=1$. The negative values of $\eta^{\theta}{}_{\theta}$ at the ridge at $\theta=1/4$ with $|\eta^{\theta}{}_{\theta}|\gg 1$ explain why the axion field experiences a tachyonic instability. It falls to the nearest of the two axion valleys, $\theta=0$ or $\theta=1/2$, and remains strongly stabilized there since $\eta^{\theta}{}_{\theta}$ at the valleys is positive and big. The same happens for all other values of $A\gtrsim 0.1$.
Using the same parameters $\beta=12^{3}$ and $\alpha=1/3$ as in our other estimates in this paper, we find that the slow-roll regime with respect to the axion field breaks down already at $A\geq 0.02$, and this breaking becomes increasingly stronger at larger values of $A$, see Fig. 3. This means that for $A\gtrsim 0.02$, the axion field experiences a strong tachyonic instability in the vicinity of the ridges at $\theta=1/4+k/2$, rapidly falls down to one of the axion valleys at $\theta=k/2$, and remains strongly stabilized there. Once the axion is stabilized, the difference between $j(\tau)$ and $\overline{j(\tau)}$ disappears, and the potential along the $\theta=1/2+k$ valleys coincides with the original potential (2). However, in this new regime, the axion field in the axion valleys is heavy, $\eta^{\theta}{}_{\theta}(0)>1$. Therefore, no isocurvature perturbations are generated in this scenario.
@@
The point $\varphi=\theta=0$ is unstable, and the fields eventually fall to one of the two minima, $\theta=\pm 0.5$, $\varphi\approx-0.101$ (or $\tau=e^{\pi i/3}$ or $e^{2\pi i/3}$). This happens because of a tachyonic instability, just as in the Higgs model or in the hybrid inflation scenario. A detailed theory of this process can be found in Felder:2000hj ; Felder:2001kt . As pointed out in Kallosh:2024ymt , in the model of the type of (2), this process is not inflationary unless $\beta$ is very close to $1$. This conclusion is valid in the model (4) near the saddle point as well. Indeed, one can check that the slow-roll parameter $\eta^{\theta}{}_{\theta}$ is large and negative. Indeed, we find that at the saddle point $\varphi=\theta=0$ in the model with $\beta=12^{3},A=0.1$ one has $\eta^{\theta}{}_{\theta}\approx-79.23$.
@@
with the universal dependence on the number of e-folds $N_{e}$.
@@ 4 Stabilized potentials in Killing coordinatess
However, the clear distinction between the (infinite number of) very sharp ridges downstairs and the very wide plateau upstairs is actually a mirage, an optical illusion created by hyperbolic geometry: in fact, as was shown in Kallosh:2024ymt ; Kallosh:2024pat , each of these ridges is physically equivalent to the entire upper part of the potential shown in Fig. 7. This means, in particular, that each of these ridges can support inflation just as well as the upper part of the half-plane shown in Fig. 7. This statement is rather nontrivial and counter-intuitive, and it raises many additional questions. For example, one may wonder what the inflationary trajectories may look like at $\varphi<0$. As we will see, investigation of the potentials with the axion valleys provides a simple answer to this question.
@@
Figure 8: This figure shows the identical scalar potentials as in Figure 7, but now in Killing coordinates (22). The left panel, 8(a), shows the potential (17) for $\gamma=0$, equivalent to (3), in Killing coordinates. The right panel, 8(b), shows the potential for $\gamma=1/4$ and is thus equal to (20). Both potentials are shown for $A=10$ and $13\alpha=$ 1, $\beta=12^{3}$. Again, on the left, valleys end in the Minkowski minima or in the dS saddle pointsn the right, all valleys end in the Minkowski minima.
@@
Thus, inflation is indeed possible at the ridge $\varphi<0$, $\theta=0$, and in fact at all other ridges shown in Fig. 8(a), see a detailed discussion of this issue in Kallosh:2024pat . Moreover, now we can visually identify all stable inflationary trajectories: they correspond to all axion valleys going down from every ridge shown in Figs.  7(a) and 8(a). These have a simple mathematical interpretation: they correspond to the lines where the function $j(\tau)$ takes real values, i.e. $j(\tau)=\overline{j(\tau)}$.
@@ Appendix A Supergravity $SL(2,\mathbb{Z})$ models with stabilized axion
If $\Lambda=|F_{X}|^{2}-3|W_{0}|^{2}>0$ and at the end of inflation $V(T,\bar{T})=0$, the total potential in eq. (29) has de Sitter vacuum. This action is $SL(2,\mathbb{Z})$ invariant if $V(T,\bar{T})$ is $SL(2,\mathbb{Z})$ invariant.

### v3

@@ 1 Introduction
We will demonstrate in this paper that one can stabilize the axion in an $SL(2,\mathbb{Z})$-compatible manner in the models based on $j$-functions Kallosh:2024ymt . Note that $j$-function and its complex conjugate $\overline{j(\tau)}$ are $SL(2,\mathbb{Z})$ invariant, so any function of $j$ and $\overline{j(\tau)}$ is $SL(2,\mathbb{Z})$ invariant. This allows to introduce axion stabilization via terms such as $|j(\tau)\pm\overline{j(\tau)}|^{2}$, or more general terms, such as $|j(\tau)\,e^{-i2\pi\gamma}-\overline{j(\tau)}\,e^{i2\pi\gamma}|^{2}$ where $\gamma$ is some constant. As we will see, this leads to the formation of axion valleys, stabilizing the axion field. Inflation in such two-field models typically occurs in the same way as in the single-field $\alpha$-attractors, and therefore leads to the same cosmological predictions.
In comparison, other models described in Kallosh:2024ymt have $SL(2,\mathbb{Z})$ invariant potentials depending on a combination of modular forms, which transform under $SL(2,\mathbb{Z})$, and on a complex field $\tau$. These involve a Dedekind function $\eta(\tau)$, a modular form of weight 1/2, an almost holomorphic modular form of weight 2, $\tilde{G}_{2}$ etc. It is not immediately clear how to construct $SL(2,\mathbb{Z})$ invariant axion stabilization terms in these models; see Kallosh:2024whb ; 6authors for an investigation of the cosmological consequences of the models without axion stabilization.
@@ 2 Axion stabilization preserving $SL(2,\mathbb{Z})$ invariance
In the present paper, we will outline a number of possibilities to avoid this potential problem by making the axion massive and stabilizing its value during inflation. As a simple example of a model with axion stabilization, consider a potential
@@
Indeed, this leads to the axion stabilization at $\theta=k/2$.
@@
Figure 2: Left panel (2(a)): A stream plot of the slow-roll inflationary trajectories for $A=10^{{-3}}$, $13\alpha=$ 1 in the range $20<\theta<1/$ 2, $65<\varphi<$ 6. Even the trajectories beginning very close to the ridge at $\theta=1/4$ gradually reach the axion valleys at $\theta=0$ or $\theta=1/2$. This process is even more efficient for the trajectories at greater values of $\varphi$ and $A$. Right panel (2(b)): The same stream plot, for $A=10^{{-2}}$.
@@
Figure 3: The slow-roll parameter $\eta^{\theta}{}_{\theta}$ (14) at $\varphi=2$ for $13\alpha=$ 1, $\beta=12^{3}$. In the left panel $A=1/10$, in the right one $A=1$. The large negative values of $\eta^{\theta}{}_{\theta}$ at the ridge at $\theta=1/4$ with $|\eta^{\theta}{}_{\theta}|\gg 1$ explain why the axion field experiences a tachyonic instability. It rapidly falls to the nearest of the two axion valleys, $\theta=0$ or $\theta=1/2$, and remains strongly stabilized there since $\eta^{\theta}{}_{\theta}$ at the valleys is positive and big. The same happens for all other values of $A\gtrsim 0.1$.
Using the same parameters $\beta=12^{3}$ and $\alpha=1/3$ as in our other estimates in this paper, we find that the slow-roll regime with respect to the axion field breaks down already at $A\geq 0.02$, and this breaking becomes increasingly stronger at larger values of $A$, see Fig. 3. This means that for $A\gtrsim 0.02$, the axion field experiences a strong tachyonic instability in the vicinity of the ridges at $\theta=1/4+k/2$, rapidly falls down to one of the axion valleys at $\theta=k/2$, and remains strongly stabilized theree, see Fig. 4.
Once the axion is stabilized, the difference between $j(\tau)$ and $\overline{j(\tau)}$ disappears, and the potential along the $\theta=1/2+k$ valleys coincides with the original potential (2). However, in this new regime, the axion field in the axion valleys is heavy, $\eta^{\theta}{}_{\theta}(0)>1$. Therefore, no isocurvature perturbations are generated in this scenario.111 Note that after the field $\theta$ falls to the axion valleys at $A\geq 0.02$ we effectively have a single field inflation along the geodisics because the field $\theta$ is heavy and stabilized at the axion valleys $\theta=k/2$. One can clearly see it in Fig. 4. This is different from multifield inflation models with non-geodesic/rapidly turning trajectories, where the “no $\eta$-problem” issue was described in Chakraborty:2019dfh ; Aragam:2021scu ; Bhattacharya:2022fze .
@@
The point $\varphi=\theta=0$ is unstable, and the fields eventually fall to one of the two minima, $\theta=\pm 0.5$, $\varphi\approx-0.101$ (or $\tau=e^{\pi i/3}$ or $e^{2\pi i/3}$). This happens because of the tachyonic instability, just as in the Higgs model or in the hybrid inflation scenario. A detailed theory of this process can be found in Felder:2000hj ; Felder:2001kt . As pointed out in Kallosh:2024ymt , in the model of the type of (2), this process is not inflationary unless $\beta$ is very close to $1$. This conclusion is valid in the model (4) near the saddle point as well. Indeed, one can check that the slow-roll parameter $\eta^{\theta}{}_{\theta}$ is large and negative. Indeed, we find that at the saddle point $\varphi=\theta=0$ in the model with $\beta=12^{3},A=0.1$ one has $\eta^{\theta}{}_{\theta}\approx-79.23$.
@@
with the universal dependence on the number of e-folds $N_{e}$. These predictions provide a good fit to the Planck/BICEP/Keck constraints on $n_{s}$ and $r$ BICEP:2021xfz .
These results are valid in the standard single-field versions of $\alpha$-attractors, where the potential at its minimum is much smaller than the height of the plateau. This condition is satisfied for $\beta=j[1]=12^{3}$ which we discussed above.
However, as mentioned in Kallosh:2024ymt , in the models with much smaller values of $\beta$, the minimum of the inflaton potential along any fixed direction except $\theta=0.5$ becomes considerably uplifted. The resulting potential shown in Fig. 6 of Kallosh:2024ymt becomes very similar to the potential of the $\alpha$-attractor versions of hybrid inflation Kallosh:2022ggf . As shown in Kallosh:2022ggf , such uplift of the $\alpha$-attractor potential results in a controllable increase of the value of $n_{s}$, which may make $SL(2,\mathbb{Z})$ models compatible with the recent ACT result $n_{s}=0.9743\pm 0.0034$ Louis:2025tst . For a recent discussion of this possibility in relation to the ACT results, see Kallosh:2025ijd .
@@ 4 Stabilized potentials in Killing coordinatess
However, the clear distinction between the (infinite number of) very sharp ridges downstairs and the very wide plateau upstairs is actually a mirage, an optical illusion created by hyperbolic geometry: in fact, as was shown in Kallosh:2024ymt ; Kallosh:2024pat , each of these ridges is physically equivalent to the entire upper part of the potential shown in Fig. 7. This means, in particular, that each of these ridges can support inflation just as well as the upper part of the half-plane shown in Fig. 7. This statement is rather non-trivial and counterintuitive, raising many additional questions. For example, one may wonder what the inflationary trajectories may look like at $\varphi<0$. As we will see, investigation of the potentials with the axion valleys provides a simple answer to this question.
@@
Figure 8: This figure shows the identical scalar potentials as in Figure 7, but now in Killing coordinates (22). The left panel, 8(a), shows the potential (17) for $\gamma=0$, equivalent to (3), in Killing coordinates. The right panel, 8(b), shows the potential for $\gamma=1/4$ and is thus equal to (20). Both potentials are shown for $A=10$ and $13\alpha=$ 1, $\beta=12^{3}$. Again, on the left, valleys end in the Minkowski minima or in the dS saddle points; on the right, all valleys end in the Minkowski minima.
@@
Thus, inflation is indeed possible at the ridge $\varphi<0$, $\theta=0$, and in fact at all other ridges shown in Fig. 8(a), see a detailed discussion of this issue in Kallosh:2024pat . Moreover, now we can visually identify all stable inflationary trajectories: they correspond to all axion valleys going down from every ridge shown in Figs.  7(a) and 8(a). These have a simple mathematical interpretation: they correspond to the lines where the function $j(\tau)$ takes real values, i.e., $j(\tau)=\overline{j(\tau)}$.
@@ Appendix A Supergravity $SL(2,\mathbb{Z})$ models with stabilized axion
If $\Lambda=|F_{X}|^{2}-3|W_{0}|^{2}>0$ and at the end of inflation $V(T,\bar{T})=0$, the total potential in eq. (29) has a de Sitter vacuum. This action is $SL(2,\mathbb{Z})$ invariant if $V(T,\bar{T})$ is $SL(2,\mathbb{Z})$ invariant.